# Electroweak Monopole-Antimonopole Pair Production at LHC

### 16 Mar 2024 v1

@@ Abstract
One of the urgent issues in high energy physics is the experimental confirmation of the electroweak monopole predicted by the standard model, and currently MoEDAL at LHC is actively searching for the monopole. However, the present LHC cannot produce the monopole if the mass is bigger than 7 TeV, while the monopole mass is expected to be around $M_{W}/\alpha\simeq 11~{}\text{TeV}$. In this paper we discuss how LHC could circumbent this energy constraint and produce the monopole even when the mass is bigger than 7 TeV, based on the following ideas. First, in the topological production of the monopole the baby monopole mass at creation could be considerably smaller than the adolescent mass. Second, the binding energy of the monopole-antimonopole pair could effectively reduce the mass of the bound state. We discuss how these ideas can actually be realized at LHC to produce the monopole pairs. In particular, we argue that LHC could produce the baby electroweak monopoles whose mass could be around 5.2 TeV, smaller than the adolescent monopole mass around 11.0 TeV. Moreover, we show that LHC could produce the monopolium bound state with mass around 5.3 TeV, even when the total mass of the monopole-antimonopole pair is around 22.0 TeV. Our analysis could play an important role for MoEDAL experiment.
@@ I Introduction
Because of these distinctive features, MoEDAL detector could identify the monopole without much difficulty, if LHC could produce it. However, the 14 TeV LHC may have no chance to produce the monopole if the mass becomes larger than 7 TeV. This is problematic, because the monopole mass coud turn out to be bigger than this, around 11.0 TeV. If this is the case, we may have little chance to detect the monopole at LHC. In this case we may have to try to detect the remnant monopoles at present universe produced in the early universe [27], or wait for the next upgrade of LHC.
However, two observations could make the monopole signal at LHC non-zero even when the mass becomes over 7 TeV. First, the monopole has to be created in pairs at LHC. This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and antimonopole, or the quantized magnetic flux of the monopole-antimonopole pair. In this case there is the possibility that the binding energy could reduce the bound state of monopole-antimonopole pair below 14 TeV, even if the monopole mass is bigger than 7 TeV.
@@
A unique feature of this topological thermal fluctuation mechanism is that this could allow LHC to produce the electroweak monopole even when the mass is heavier than 7 TeV. To understand this point we first have to keep in mind that the monopole mass we have been talking about is the mass at the zero temperature. However, in this thermal production the monopole mass depends on the temperature, so that at the creation the initial monopole mass could be smaller than the zero temperature mass [27]. This tells that we have to distinguish the finite temperature monopole mass from the zero temperature monopole mass. So, from now on we call the initial monopole mass (the monopole mass at birth) as the “baby” monopole mass and the zero temperature mass as the “adolescent” monopole mass.
The purpose of this paper is to discuss how LHC can accommodate the above ideas and produce the electroweak monopoles even when the monopole mass exceeds 7 TeV. We show that in the topological thermal production the monopole production at LHC starts from $T_{c}\simeq 147~{}\text{TeV}$ and stops at $T_{min}^{LHC}\simeq 119~{}\text{TeV}$ before the Ginzburg temperature $T_{G}\simeq 57.5~{}\text{GeV}$, with the monopole mass around 5.3 TeV. Moreover, we show that LHC could produce the monopolium bound states of mass around 5.3 TeV, even though the mass of the monopole-antimonopole pair becomes 22.0 TeV.
As we will see, just after the phase transition the baby monopole mass could be as low as 1.0 TeV. Notice that this would have been impossible if LHC produces the monopole by Drell-Yan or Schwinger mechanism, where the concept of the temperature dependent monople mass does not exist. On the other hand, we emphasize that the monopole production in the early universe and the monopole production at LHC have important differences. First, the monopole production in the early universe stops around the Ginzburg temperature, while the monopole production at LHC stops at higher temperature. This is because at the present LHC the monopole mass can not exceed 7 TeV. So LHC has less monopole production time. Second, the size of the thermal radiation bath in the early universe is limited only by the horizon distance, but the size of the fireball at LHC in p-p collision is fixed by the proton size which is very small. Because of this, the monopole in the early universe has unlimited energy nearby but the monopole at LHC has very limited energy coming from the p-p fireball. These differences play important roles in the monopole production at LHC.
@@
This makes the monopolium bound state another probable signal for the monopole production at LHC. The monopolium itself has been studied before for various reasons [32, 33, 34]. For example, Nambu proposed to view the monopolium as a model of strong interaction, because the magnetic coupling of the monopole-antionopole pair drastically increases the energy spectrum to the order of hundred MeV, the mass scale of the low lying hadrons [34]. But the electroweak monopolium made of the Cho-Maison monopole pair has not been studied before.
Our study on the electroweak monopolium tells the followings. First, the naive atomic model of monopolium does not describe a realistic monopolium, because it has unrealistically big binding energy and too small size. This is because the magnetic Coulombic coupling of the monopole-antimonopole pair is too strong and the monopole mass is huge.
@@ II Electroweak Monopole Production Mechanisms at LHC
We emphasize that this is not the Weinberg-Salam Lagrangian in the unitary gauge. This is the gauge independent abelianization of the Weinberg-Salam Lagrangian. Notice that in this abelianization the Higgs doublet disappears completely, and the Higgs, W, and Z boson masses are given by $M_{H}=\sqrt{\lambda}\rho_{0}$, $M_{W}=g\rho_{0}/2$, $M_{Z}=\sqrt{g^{2}+g^{\prime 2}}\rho_{0}/2$.
The Lagrangian has the monopole topology $\pi_{2}(S^{2})$ that comes from the unit doublet $\xi$ which could be identified as a $\mathbb{C}P^{1}$ field. It has monopole solutions, the naked Cho-Maison monopole and the Cho-Maison monopole dressed by the W and Higgs bosons [14]. Although the Cho-Maison monopole has infinite energy classically, one could predict the mass to be around 4 to 11 TeV. Intuitively we could argue the mass to be $1/\alpha$ times heavier than the W boson mass, around 11 TeV. This is because the monopole mass comes from the same Higgs mechanism that makes the W boson massive, except that here the gauge coupling is magnetic (i.e., $4\pi/e$). We could backup this arguement regularizing the Cho-Maison monopole, and estimate the mass to be around 4 to 10 TeV [23, 24, 25, 27, 28]. In the following we will assume for simplicity that the electroweak monopole mass is $M_{W}/\alpha$, around 11.0 TeV.
@@
To understand the electroweak monopole production at LHC we have to know the monopole production mechanism. There are two contending monopole production mechanisms in particle physics community at present, the Drell-Yan (and two photon fusion) process and the Schwinger mechanism. In the Drell-Yan process the monopole-antimonopole pairs are produced by the electromagnetic interaction through the photon [19, 20]. But in the Schwinger mechanism a strong magnetic field creates the monopole-antimonopole pairs, just like a strong electric background creates the electron-positron pairs in the non-perturbative QED [22, 29, 30]. The justification for this comes from the fundamental assumption that a theory of monopole must have the dual symmetry, i.e., must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ with $(e,g)$ to $(g,-e)$. The Drell-Yan process is shown graphically in Fig. 1.
@@
A unique feature of this mechanism is that the monopole production is not controlled by any fundamental interaction, but by thermal fluctuation which does not involve any fundamental constant. Perhaps more importantly, in this thermal fluctuation the monopole mass at creation (the baby monopole mass) could be considerably smaller than the adolescent monopole mass. This is because in this thermal production, the monopole mass depends on the temperature at creation. As we will see in the following this could have a deep impact on the monopole production at LHC.
@@ III Monopole Production in Early Universe: A Review
We can find it graphically from (21) and (28). This is shown in Fig. 4. From this we have [27]
@@
The effective potential at the Ginzburg temperature is shown in Fig. 3. Notice that at $\rho_{+}(T_{G})$ is quite close to the Higgs vacuum at zero temperature, which confirms that the electroweak monopole production lasts long time after the phase transition.
@@
This tells that at $T_{G}$ already becomes close to the Higgs mass at zero temperature. But the W-boson is massless at high temperature above $T_{c}$ (before the symmetry breaking), and becomes massive toward the value 7.1 GeV at $T_{c}$ as the Higgs vacuum $\rho=0$ tunnels to $\rho_{+}(T_{c})\simeq 21.7~{}\text{GeV}$. Moreover, we have
@@
This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~{}\text{TeV}$), so that at $T_{G}$ the monopole mass acquires almost the adolecent value. In Table I we list the masses of Higgs, W boson, and electrowak monopole at different temperatures for comparison.
The unique feature of the topological monopole production mechanism is that (unlike the Drell-Yan or Schwinger mechanism) the monopole mass depends on the temperature at creation, and that the baby monopole mass could be considerably smaller than the adolescent monopole mass. This could allow LHC to produce the monopole even when the adolescent mass is bigger than 7 TeV.
@@ IV Electroweak monopoles from a fireball at LHC
To do that, we estimate the temperature at which the monopole production stops at the present LHC. Since LHC should produce the monopole in pairs, the minimum temperature $T_{min}$ at the 14 TeV LHC for the monopole production must satisfy the energy condition
$M_{max}=7~{}\text{TeV}\simeq\frac{\bar{M}_{W}(T_{min})}{\alpha}=\frac{g}{2}\frac{\rho_{+}(T_{min})}{\alpha}.$ (41)
@@
If so, at LHC the electroweak monopole production starts at $T_{c}$ around 146.7 GeV and stops at $T_{min}^{LHC}$ around 119.4 GeV, not at the Ginzburg temperature. Moreover, in average the monopole production temperature at LHC is given by
@@
not at $T_{i}$ given by (30).
This tells that LHC starts to produce the electroweak monopole with mass 1.0 TeV at around 146.7 GeV, and stops producing the monopole with mass 7 TeV at 119.4 TeV. In average LHC produces the infant electroweak monopole mass around 5.3 TeV, much less than the adolescent mass 11.0 TeV. This implies that the 14 TeV LHC could actually produce the electroweak monopole pair even when the mass of the monopole pair becomes bigger than 14 TeV. This is remarkable.
@@
However, some of the infant monopole-antimonopole pairs at LHC could settle down to form a monopolium bound state before they escape from the fireball. This necessiate us to study the monopolium bound state which we discuss in the following.
@@ V Electroweak Monopolium: Atomic Model
To discuss the atomic model of the monopolium we let the monopole mass be $M$ and consider the Schrödinger equation of the monopolium wave function
@@
where $\mu=M/2$ is the reduced mass and $V$ is the Coulombic magnetic potential of the monopole-antimonopole pair. Obviously this becomes formally identical to the Schrödinger equation of the Hydrogen atom, except that the coupling strenth of the potential is replaced by $4\pi/e$ and the electron mass is replaced by $\mu$.
@@
$\simeq-\frac{50,207}{n^{2}}~{}{\rm TeV},$ (50)
where $\alpha=e^{2}/4\pi$ is the fine structure constant and we have put $M=M_{W}/\alpha\simeq 10.7~{}\text{TeV}$. So the Rydburg energy of the monopolium becomes 50,207 TeV. Moreover, the Bohr radius $R$ of the monopolium is given by
@@
The reason for the above results originates from two facts. First, the magnetic coupling of the monopole is too strong. Second, the monopole mass is too big. So we could make the atomic model more realistic by improving on these points. For example, we could modify the magnetic Coulombic potential at short distances for $r\leq 1/M_{W}$, to accommodate the fact that the monopole has a finite size $1/M_{W}$. Moreover, we could make the monopole mass much less than 11 TeV because the monopole, at the initial stage of the formation, has a temperature dependent mass that is much less than the adolescent mass.
@@ VI Electroweak Monopolium: A Realistic Model
Notice that the above equation determines the energy spectrum of the monopolium, so that it is very important to solve (8). Unfortunately, this condition can only be solved numerically, and we display the energy spectrum in Fig. 8 (for $l=0$).
@@
$\simeq-16.3~{}\text{TeV}.$ (84)
where we have put $M\simeq 10.7~{}\text{TeV}$. Compared with (50), this is roughly $3\times 10^{-4}$ times reduction of the naive estimate of the binding energy given by hydrogen-like atomic model of the monopolium. In particular, this tells that the monopole-antimonopole pair with mass 21.4 TeV makes the monopolium bound state of mass of 5.1 TeV at the ground state. This is precisely what we have hoped for.

### 19 Mar 2024 v2

@@ Abstract
One of the urgent issues in high energy physics is the experimental confirmation of the electroweak monopole predicted by the standard model, and currently MoEDAL at LHC is actively searching for the monopole. However, the present LHC cannot produce the monopole if the mass is bigger than 7 TeV, while the monopole mass is expected to be around $M_{W}/\alpha\simeq 11~{}\text{TeV}$. In this paper we discuss how LHC could circumbent this energy constraint and produce the monopole even when the mass becomes bigger than 7 TeV, based on the following ideas. First, in the topological production of the monopole the baby monopole mass at creation could be considerably smaller than the adolescent mass. Second, the binding energy of the monopole-antimonopole pair could effectively reduce the mass of the bound state. We discuss how these ideas can actually be realized at LHC to produce the monopole pairs. In particular, we argue that LHC could produce the baby electroweak monopoles whose mass could be around 5.3 TeV, smaller than the adolescent monopole mass around 11.0 TeV. Moreover, we show that LHC could produce the monopolium bound state with mass around 2.5 TeV, even when the total mass of the baby monopole-antimonopole pair becomes around 10.6 TeV. Our analysis could play an important role for MoEDAL experiment.
@@ I Introduction
Because of these distinctive features, MoEDAL detector could identify the monopole without much difficulty, if LHC could produce it. However, the 14 TeV LHC may have no chance to produce the monopole if the mass becomes larger than 7 TeV. This is problematic, because the monopole mass could turn out to be bigger than this, around 11.0 TeV. If this is the case, we may have little chance to detect the monopole at LHC. In this case we may have to try to detect the remnant monopoles at present universe produced in the early universe [27], or wait for the next upgrade of LHC.
However, two observations could make the monopole signal at LHC non-zero even when the mass becomes over 7 TeV. First, the monopole has to be created in pairs at LHC. This implies that at the initial stage, the monopoles could appear in the form of the atomic monopolium states made of the monopole and antimonopole, or the quantized magnetic flux of the monopole-antimonopole pair. In this case there is the possibility that the binding energy could reduce the bound state of monopole-antimonopole pair below 14 TeV, even if the monopole mass becomes bigger than 7 TeV.
@@
A unique feature of this topological thermal fluctuation mechanism is that this could allow LHC to produce the electroweak monopole even when the mass becomes heavier than 7 TeV. To understand this point we first have to keep in mind that the monopole mass we have been talking about is the mass at the zero temperature. However, in this thermal production the monopole mass depends on the temperature, so that at the creation the initial monopole mass could be smaller than the zero temperature mass [27]. This tells that we have to distinguish the finite temperature monopole mass from the zero temperature monopole mass. So, from now on we will call the initial monopole mass (the monopole mass at birth) as the “baby” monopole mass and the zero temperature mass as the “adolescent” monopole mass.
[…]
As we will see, just after the phase transition the baby monopole mass could be as low as 1.0 TeV. This would have been impossible if LHC produces the monopole by Drell-Yan or Schwinger mechanism, where the concept of the temperature dependent monople mass does not exist. On the other hand, we emphasize that the monopole production in the early universe and the monopole production at LHC have important differences. First, the monopole production in the early universe stops around the Ginzburg temperature, while the monopole production at LHC stops at higher temperature. This is because at the present LHC the monopole mass can not exceed 7 TeV. So LHC has less monopole production time. Second, the size of the thermal radiation bath in the early universe is limited only by the horizon distance, but the size of the fireball at LHC in p-p collision is fixed by the proton size which is very small. Because of this, the monopole in the early universe has unlimited energy nearby but the monopole at LHC has very limited energy coming from the p-p fireball. These differences play important roles in the monopole production at LHC.
@@
This makes the monopolium bound state a more probable signal for the monopole production at LHC. The monopolium itself has been studied before for various reasons [32, 33, 34]. For example, Nambu proposed to view the monopolium as a model of strong interaction, because the magnetic coupling of the monopole-antionopole pair drastically increases the energy spectrum to the order of hundred MeV, the mass scale of the low lying hadrons [34]. But the electroweak monopolium made of the Cho-Maison monopole pair has not been studied before.
Our study on the electroweak monopolium tells the followings. First, the naive atomic model of monopolium does not describe a realistic monopolium, because it has unrealistically big binding energy and too small size. This is because the magnetic Coulombic coupling of the monopole-antimonopole pair is too strong and the monopole mass is too big.
@@ II Electroweak Monopole Production Mechanisms at LHC
We emphasize that this is not the Weinberg-Salam Lagrangian in the unitary gauge. This is the gauge independent abelianization of the Weinberg-Salam Lagrangian. Notice that in this abelianization the Higgs doublet disappears completely, and the Higgs, W, and Z boson acquire the mass $M_{H}=\sqrt{\lambda}\rho_{0}$, $M_{W}=g\rho_{0}/2$, $M_{Z}=\sqrt{g^{2}+g^{\prime 2}}\rho_{0}/2$.
The Lagrangian has the monopole topology $\pi_{2}(S^{2})$ that comes from the unit doublet $\xi$ which could be identified as a $\mathbb{C}P^{1}$ field. It has monopole solutions, the naked Cho-Maison monopole and the Cho-Maison monopole dressed by the W and Higgs bosons [14]. Although the Cho-Maison monopole has infinite energy classically, one could predict the mass to be around 4 to 11 TeV. Intuitively we could argue the mass to be $1/\alpha$ times heavier than the W boson mass, around 11 TeV. This is because the monopole mass comes from the same Higgs mechanism that makes the W boson massive, except that here the gauge coupling is magnetic (i.e., $4\pi/e$). We could backup this arguement regularizing the Cho-Maison monopole, and estimate the mass to be around 4 to 10 TeV [23, 24, 25, 27, 28]. But in the following we will assume for simplicity that the electroweak monopole mass is $M_{W}/\alpha$, around 11.0 TeV.
@@
To understand the electroweak monopole production at LHC we have to know the monopole production mechanism. In the Drell-Yan process the monopole-antimonopole pairs are produced by the electromagnetic interaction through the photon [19, 20]. But in the Schwinger mechanism a strong magnetic field creates the monopole-antimonopole pairs, just like a strong electric background creates the electron-positron pairs in the non-perturbative QED [22, 29, 30]. The justification for this comes from the assumption that a theory of monopole must have the dual symmetry, i.e., must be symmetric under the dual transformation $(\vec{E},\vec{H})$ to $(\vec{H},-\vec{E})$ together with $(e,g)$ to $(g,-e)$. The Drell-Yan process is shown graphically in Fig. 1.
@@
A remarkable feature of this mechanism is that the monopole production is not controlled by any fundamental interaction, but by thermal fluctuation which does not involve any fundamental constant. Perhaps more importantly, in this thermal fluctuation the monopole mass at creation (the baby monopole mass) could be considerably smaller than the adolescent monopole mass. This is because in this thermal production, the monopole mass depends on the temperature at creation. As we will see in the following this could have a deep impact on the monopole production at LHC.
@@ III Monopole Production in Early Universe: A Review
We can find $T_{G}$ graphically from (21) and (28). This is shown in Fig. 4. From this we have [27]
@@
The effective potential at the Ginzburg temperature is shown in Fig. 3. Notice that $\rho_{+}(T_{G})$ is quite close to the Higgs vacuum at zero temperature, which confirms that the electroweak monopole production lasts long time after the phase transition.
@@
This tells that at $T_{G}$ it already becomes close to the Higgs mass at zero temperature. But the W-boson is massless at high temperature above $T_{c}$ (before the symmetry breaking), and becomes massive toward the value 7.1 GeV at $T_{c}$ as the Higgs vacuum $\rho=0$ tunnels to $\rho_{+}(T_{c})\simeq 21.7~{}\text{GeV}$. Moreover, we have
@@
This implies that the infant monopole masses at $T_{c}$ and $T_{G}$ are (with $\bar{M}\simeq\bar{M}_{W}/\alpha$) around 1.0 TeV and 10.4 TeV (assuming the adolescent mass is $M_{W}/\alpha\simeq 11.0~{}\text{TeV}$). This tells that the baby monopole mass can be considerably smaller than the adolescent monopole mass, although at $T_{G}$ the monopole mass becomes close to the adolecent value. This could allow LHC to produce the monopole even when the adolescent mass becomes bigger than 7 TeV. In Table I we list the masses of Higgs, W boson, and electrowak monopole at different temperatures for comparison.
@@
At this point one might wonder if we could also observe the Higgs and W bosons which have the temperature dependent masses (the “baby” Higgs and W bosons) at LHC. This is not possible, because they are ordinary particles produced by the ordinary process, not the topological particles produced topologically. So there is no reason to believe that LHC can produce them.
@@ IV Electroweak monopoles from a fireball at LHC
To do that, we estimate the temperature at which the monopole production stops at the present LHC. Since LHC should produce the monopole in pairs, the minimum temperature $T_{min}^{LHC}$ at the 14 TeV LHC for the monopole production must satisfy the energy condition
$M_{max}=7~{}\text{TeV}\simeq\frac{\bar{M}_{W}(T_{min}^{LHC})}{\alpha}=\frac{g}{2}\frac{\rho_{+}(T_{min}^{LHC})}{\alpha}.$ (41)
@@
If so, at LHC the electroweak monopole production starts at $T_{c}$ around 146.7 GeV and stops at $T_{min}^{LHC}$ around 119.4 GeV, not at the Ginzburg temperature. In average the monopole production temperature at LHC is given by
@@
not at $T_{i}$ given by (30). This tells that LHC starts to produce the electroweak monopole with mass 1.0 TeV at around 146.7 GeV, and stops producing the monopole with mass 7 TeV at 119.4 TeV. In average LHC produces the infant electroweak monopole mass around 5.3 TeV, much less than the adolescent mass 11.0 TeV. This implies that the 14 TeV LHC could actually produce the electroweak monopole pair even when the mass of the monopole pair becomes bigger than 14 TeV. This is remarkable.
@@
On the other hand, some of the infant monopole-antimonopole pairs at LHC could settle down to form a monopolium bound state before they escape from the fireball. This necessiate us to study the monopolium bound state which we discuss in the following.
@@ V Electroweak Monopolium: Atomic Model
Let the monopole mass be $M$ and consider the Schrödinger equation of the monopolium wave function
@@
where $\mu=M/2$ is the reduced mass and $V$ is the Coulombic magnetic potential of the monopole-antimonopole pair. Obviously this is formally identical to the Schrödinger equation of the Hydrogen atom, except that the coupling strenth of the potential is replaced by $4\pi/e$ and the electron mass is replaced by $\mu$.
@@
$\simeq-\frac{51,615.3}{n^{2}}~{}{\rm TeV},$ (50)
where $\alpha=e^{2}/4\pi$ is the fine structure constant and we have put $M=M_{W}/\alpha\simeq 11.0~{}\text{TeV}$. So the Rydburg energy of the monopolium becomes 51.615 TeV. Moreover, the Bohr radius $R$ of the monopolium is given by
@@
The reason for the above results originates from two facts. First, the magnetic coupling of the monopole is too strong. Second, the monopole mass is too big. We could make the atomic model more realistic by improving on these points. For example, we could modify the magnetic Coulombic potential at short distances for $r\leq 1/M_{W}$, to accommodate the fact that the monopole has a finite size $1/M_{W}$. Moreover, we could make the monopole mass less than 11 TeV because the monopole, at the initial stage of the formation, has a temperature dependent mass that is much less than the adolescent mass.
@@ VI Electroweak Monopolium: A Realistic Model
Notice that the above equation determines the energy spectrum of the monopolium, so that it is very important to solve (80). Unfortunately, this condition can only be solved numerically, and we display the energy spectrum in Fig. 8 for $l=0$.
@@
$\simeq-16.8~{}\text{TeV}.$ (84)
where we have put $M\simeq 11.0~{}\text{TeV}$. Compared with (50), this is roughly $3\times 10^{-4}$ times reduction of the naive estimate of the binding energy given by hydrogen-like atomic model of the monopolium. In particular, this tells that the monopole-antimonopole pair with mass 22.0 TeV makes the monopolium bound state of mass of 5.2 TeV at the ground state. This is precisely what we have hoped for.
@@
Notice that in the above analysis we have assumed that the monopole mass is 11 TeV. However, at LHC the monopole mass at $T_{min}^{LHC}$ becomes 5.3 TeV, so that the monopolium at LHC is expected to be the monopole-antimonopole bound state with mass 10.6 TeV. This will change the energy spectrum and the Bohr radius of the monopolium. For example, $E_{0}$ in (84) should change to $E_{0}\simeq-8.1~{}\text{TeV}$, which reduces the monopolium mass to 2.5 TeV at LHC.